# Anomaly Detection in Complex Dynamical Systems: A Systematic Framework Using Embedding Theory and Physics-Inspired Consistency

### v2

@@ III-B4 Consistency Metrics
III-C Anomaly Detection Logic
We integrate the developed mathematical foundations into anomaly detection by leveraging state-derivative pairs in the latent space. Normal states are assumed to lie within a confined region, with deviations indicating anomalies.
Training uses the first 60% of each engine’s data, with an 80/20 split into training and validation sets. Thresholds for each latent node are set as percentiles of the validation data and optimized using a classification metric (e.g., F1-score). This optimization can be performed based on the classification performance on the validation dataset. The threshold remains fixed across all engines. During testing, the model infers the derivative representation from a single time step, as measurements at time $t$ inherently encode rate changes.
Detection logic:
  • If all latent nodes exceed their thresholds during normal operation, it is classified as a false positive (FP); otherwise, it is a true negative (TN).
  • If at least one latent node exceeds its threshold in the anomaly range, it is classified as a true positive (TP); otherwise, it is a false negative (FN).
Detection rules can be adapted based on specific use cases and risk strategies, such as requiring multiple nodes to exceed thresholds before flagging an anomaly.
@@ IV-B Phase Space Representation and State-Derivative Relationships
Figure 2: (a) Phase-space plot of the latent node pair (2,6) average. (b) State-derivative consistency metric for this pair. (c) A detailed view of (a) (highlighted grey box), showing phase-space trajectories and the attractor region at the beginning of the system dynamics. The timestep color map applies to all figures.
Fig. 3a-f presents two state-derivative pairs along with their respective phase space representations. A gradual increase or decrease in both state and derivative nodes is observed as the system transitions toward the anomalous range. In the phase space plots, we applied a sliding window of three and slightly smoothed the node values for visualization purposes. This processing was not used in the actual anomaly detection model. During the initial phase of the system dynamics, the states evolve more gradually, supporting the hypothesis that, in the normal regime, the dynamics follow a simple and stable attractor geometry. The phase space representation further indicates that normal states are embedded more densely than anomalies, with a continuous trend toward the anomalous region.
@@
Let’s have a more detailed view on the phase space evolution. Fig. 2a illustrates a clear drift toward the anomalous region, highlighting a gradual deviation in system behavior. Fig. 2b presents the evolution of $\rho$, which quantifies the causal approximation of state-derivative pairs. In the normal range, $\rho$ remains stable on average, indicating a consistent dynamical description. However, as the system enters the anomalous regime, a significant increase in $\rho$ is observed, suggesting a loss in the dynamical consistency of the system. The theoretical derivation showed that when system dynamics deviate from a stable state, the increasing complexity requires a higher-dimensional embedding to adequately capture the system’s behavior. We hypothesized that this effect would lead to the collapse of state-derivative approximations, serving as an indicator of anomalous behavior. Our observations now confirm this, as we see a loss of dynamical consistency in the latent space, reinforcing our hypothesis that anomalies emerge when the system transitions beyond the representational capacity of the learned embedding.
In Fig. 2c, we observe the evolution of the phase space representation over time, transitioning from normal to abnormal behavior. Initially, the phase space exhibits an attractor-like structure around which the dynamics evolve in a stable manner. As the system progresses, we observe increased dispersion and a gradual drift away from the attractor toward the anomalous region. While in the range characterized as normal, the state-derivative approximation remains stable, as seen in Fig. 2a. However, subtle deviations from the attractor’s dynamics can already be detected in the earliest phase of the transition. This aligns with our second hypothesized effect, which suggests that even minor disruptions in the system’s stability could serve as an early indicator of anomalies.
These findings reinforce that our model is not merely a data-driven classifier for anomaly detection but instead captures meaningful insights into the underlying system dynamics.
@@ IV-C Comparison with Literature Benchmarks
TABLE II: Performance comparison of different approaches benchmarks for anomaly detection in the C-MAPSS dataset.
Algorithm Acc. (%) Prec. (%) Rec. (%) F1 (%)
TDC-AE $\mathbf{99.3}\pm 0.3$ $\mathbf{99.99}\pm 0.03$ $\mathbf{98.30}\pm 0.07$ $\mathbf{99.14}\pm 0.04$
DeepLSTM-AE [19] 96.45 94.81 98.12 96.44
Random Forest [19] 92.33 97.40 86.66 91.71
CNN-AE [19] 92.22 89.08 95.86 92.35
SMOTE Transformer [7] 89.59 93.95 91.77 -
XGBoost [19] 84.57 77.73 96.00 85.90
Dense-AE [18] - 89.6 72.4 80.1
@@
Tab. II presents a comparison across different architectures. Our approach outperforms all reported models across all detection metrics, demonstrating consistent results across repeated attempts. The latent node threshold in all attempts was set using the 75th percentile of the validation set.
Even complex models, such as LSTM-AEs, Transformer with attention mechanism designed to capture long-range temporal correlations, do not achieve superior performance. In [19], the exact LSTM architecture is not explicitly provided, so we base our assumptions on similar studies of complex physical systems [9], considering the following configuration: sequence length $L=48$, hidden dimension $d_{\text{hidden}}=16$, and input dimension $d_{\text{input}}=24$. This results in a Multiply-Accumulate Operations (MAC) count of 245,760.
In contrast, our approach achieves superior detection performance with significantly lower computational complexity, requiring only 2,688 MACs.
@@ V Conclusion
The results demonstrate that our chosen embedding strategy can effectively capture time correlation from a single time step. TDC-AE outperforms benchmarks on the C-MAPSS dataset, surpassing computationally intensive models like LSTMs and Transformers, while achieving a nearly 200x reduction in MAC operations, keeping them below 3,000 MACs. This efficiency makes our approach particularly suitable for lightweight edge computing.
@@ A-A General Definition of an Oscillatory Dynamical System
Figure 4: A driven damped pendulum with angular displacement $\theta$, velocity $\dot{\theta}$, damping $-\gamma\dot{\theta}$, and external driving force $F$.
@@ B-B TDC-AE
[…]
@@ B-C Implementation of the TDC-AE and Hardware
TABLE III: Hyperparameters used in training the autoencoder for the C-MAPSS (FD001) dataset.
@@
The model was implemented using PyTorch 2.3.0 with CUDA 11.8, leveraging GPU acceleration to enhance computational efficiency and reduce training time. The autoencoder’s architecture and training setup are detailed in Table III.
@@ Appendix C Figures
Figure 7: Schematic of temporal difference consistency in the latent space of an autoencoder. The derivative node $z_{\dot{}}$ is aligned with the central difference of the static node $z$ using time steps $t-1$, $t$, and $t+1$.

### v3

@@ III-C4 Consistency Metrics
III-D Anomaly Detection Logic
We integrate the developed mathematical foundations into anomaly detection by leveraging state-derivative pairs in the latent space. Normal states are assumed to lie within a confined region, with deviations indicating anomalies. Training uses the first 60% of each engine’s data, with an 90/10 split into training and validation sets.
Thresholds are computed as percentiles of the training data across all engines and a moving average is applied to all latent node values. Both are optimized for F1-score on anomalous last 40% of the training dataset. During testing, the model infers the derivative representation from a single time step, leveraging the fact that measurements at time $t$ inherently capture rate changes.
Detection logic: Each latent node’s normal operating range is not predefined but dynamically recalibrated. Initially, the baseline is set using the first observed value. It is then updated using a moving average of the most recent 10 values, providing a real-time estimate of the normal distribution. An instance is classified as anomalous (positive) if at least two latent nodes simultaneously exceed their respective upper thresholds or fall below their lower thresholds; otherwise, it is considered normal (negative). Classification outcomes follow the standard notation: true negative (TN), false positive (FP), true positive (TP), and false negative (FN). In high-stakes applications like aviation, failing to raise any alarm near end-of-life is unacceptable. In addition to conventional classification metrics, we introduce the Critical Detection Rate (CDR), defined as the percentage of engines for which the detector produces at least one positive signal in the final 10% of their lifetime. Although progressive degradation can lead to early false negatives that lower recall, these are less critical than missed detections near failure.
@@ IV-B Phase Space Representation and State-Derivative Relationships
Figure 2: (a) A detailed view of (b) (highlighted grey box), showing phase-space trajectories and the attractor region at the beginning of the system dynamics. (b) Phase-space plot of the latent node pair (2, 6) average. (c) State-Derivative Consistency metric for this pair. The timestep color map applies to all figures.
Fig. 3a-f presents two state-derivative pairs along with their respective phase space representations. A gradual increase or decrease in both state and derivative nodes is observed as the system transitions toward the anomalous range. A gradual increase or decrease in both state and derivative nodes is observed as the system transitions toward the anomalous range. For the phase space plots, we slightly smoothed the node values using a moving average with a window of three. During the initial phase of the system dynamics, the states evolve more gradually, supporting the hypothesis that, in the normal regime, the dynamics follow a simple and stable attractor geometry. The phase space representation further indicates that normal states are embedded more densely than anomalies, with a continuous trend toward the anomalous region.
@@
Let’s have a more detailed view on the phase space evolution. Fig. 2a illustrates a clear drift toward the anomalous region, highlighting a gradual deviation in system behavior. Fig. 2c presents the evolution of $\rho$, which quantifies the causal approximation of state-derivative pairs. In the normal range, $\rho$ remains stable on average, indicating a consistent dynamical description. However, as the system enters the anomalous regime, a significant increase in $\rho$ is observed, suggesting a loss in the dynamical consistency of the system. The theoretical derivation showed that when system dynamics deviate from a stable state, the increasing complexity requires a higher-dimensional embedding to adequately capture the system’s behavior. We hypothesized that this effect would lead to the collapse of state-derivative approximations, serving as an indicator of anomalous behavior. Our observations now confirm this, as we see a loss of dynamical consistency in the latent space, reinforcing our hypothesis that anomalies emerge when the system transitions beyond the representational capacity of the learned embedding.
In Fig. 2b, we observe the evolution of the phase space representation over time, transitioning from normal to abnormal behavior. Initially, the phase space exhibits an attractor-like structure around which the dynamics evolve in a stable manner. As the system progresses, we observe increased dispersion and a gradual drift away from the attractor toward the anomalous region. While in the range characterized as normal, the state-derivative approximation remains stable, as seen in Fig. 2a. However, subtle deviations from the attractor’s dynamics can already be detected in the earliest phase of the transition. This aligns with our second hypothesized effect, which suggests that even minor disruptions in the system’s stability could serve as an early indicator of anomalies.
We numerically assess the injectivity of the encoder by evaluating pairwise distance ratios in latent space across more than 4,000 samples per subset. No instances were found that violated the injectivity criterion (i.e., minimum ratio below $10^{-5}$), indicating the absence of latent-space collisions.
Table II summarizes the performance of our model on the FD001 and FD003 subsets. We observe high overall accuracy, precision, recall, and F1-scores on both datasets, indicating consistent detection performance. The precision values are notably strong, and most false positives (FPs) occur near the boundary between the training (first 60% of lifetime) and testing (last 40%) regions, where the engine behavior starts to change. Similarly, the few false negatives (FNs) we observed are also concentrated around this transition region. Importantly, the Critical Detection Rate (CDR), the percentage of engines with at least one positive detection in the final 10% of their lifetime, reaches 100% on both datasets, demonstrating that the model successfully detects every failing engine before its end of life.
Please refer to Tab. IV in Appendix B-C for a detailed overview of the model architecture, training hyperparameters, and the threshold settings used for anomaly detection.
TABLE II: Performance of TDC-AE across C-MAPSS subsets
Dataset Acc. (%) Prec. (%) Rec. (%) F1 (%) CDR (%)
FD001 96.49 95.95 95.18 95.56 100.0
FD003 94.84 93.81 93.16 93.49 100.0
@@ IV-C Comparison with Literature Benchmarks
TABLE III: Performance comparison of different approaches benchmarks for anomaly detection in the C-MAPSS dataset FD001.
Algorithm Acc. (%) Prec. (%) Rec. (%) Spec.(%) F1 (%)
TDC-AE 96.49 95.95 95.18 97.35 95.56
DeepLSTM-AE [19] 96.45 94.81 98.12 - 96.44
Random Forest [19] 92.33 97.40 86.66 - 91.71
CNN-AE [19] 92.22 89.08 95.86 - 92.35
SMOTE Transformer [7] - - 92.22 93.44 -
XGBoost [19] 84.57 77.73 96.00 - 85.90
Dense-AE [18] - 89.6 72.4 - 80.1
@@
Tab. III presents a comparison across different architectures. Our approach achieves competitive results across all detection metrics. While more complex models such as LSTM-AEs and Transformers are designed to capture long-range temporal dependencies, our method attains comparable performance with a substantially simpler architecture. In [19], the exact LSTM architecture is not explicitly provided, so we base our assumptions on similar studies of complex physical systems [9], considering the following configuration: sequence length $L=48$, hidden dimension $d_{\text{hidden}}=16$, and input dimension $d_{\text{input}}=24$. This results in a Multiply-Accumulate Operations (MAC) count of 245,760.
In contrast, our approach achieves superior detection performance with significantly lower computational complexity, requiring only about 3360 MACs.
@@ V Conclusion
The results demonstrate that our embedding strategy can drastically reduce model size. On the C-MAPSS dataset, TDC-AE matches benchmark performance while achieving nearly a 100× reduction in MAC operations compared to LSTM models, bringing the computation down to about 3,000 MACs. This efficiency makes our approach particularly well-suited for lightweight edge computing.
As a next step, we plan to investigate the theoretical implications of having access to a system-level state representation, effectively a phase plane of complex systems. Exploring this higher-level view could provide new insights into the underlying dynamics, stability properties, and fault evolution mechanisms, enabling more principled approaches to monitoring and control
@@ A-A General Definition of an Oscillatory Dynamical System
Figure 4: A driven damped pendulum with angular displacement $\theta$, velocity $\dot{\theta}$, damping $\gamma\dot{\theta}$, and external driving force $F$.
@@
Definition 1 (Box-Counting Dimension [35]).
The box-counting dimension of $A$ is defined as
provided the limit exists.
Theorem 1 (Fractal Whitney Embedding Prevalence Theorem [21]).
  1. 1.
    $F$ is one-to-one on $A$.
  2. 2.
@@ B-B TDC-AE
8:   Compute reconstruction loss using MSE: $\text{Rec-Loss}\leftarrow\text{MSE}(X_{\text{rec}},X_{t})$
9:   Backpropagate the total loss: Compute gradients w.r.t. model parameters for $\text{Rec-Loss}+\alpha\cdot\text{TDC-Loss}$
@@ B-C Implementation of the TDC-AE and Hardware
TABLE IV: Hyperparameters used in training the autoencoder and derived thresholds for the C-MAPSS (FD001 and FD003) datasets.
@@
Threshold (FD001) max = 86th perc., min = 9th perc.
Threshold (FD003) max = 75th perc., min = 22nd perc.
Moving Average Window 12
@@
The model was implemented using PyTorch 2.3.0 with CUDA 11.8, leveraging GPU acceleration to enhance computational efficiency and reduce training time. The autoencoder’s architecture and training setup are detailed in Table IV.
@@ Appendix C Figures
Figure 7: Schematic of temporal difference consistency in the latent space of an autoencoder. The derivative node $z_{\dot{}}$ is aligned with the central difference of the static node $z$ using time steps t1t-1, $t$, and $t+1$.